OpenStudy (anonymous):

simplify 5x^{3} y^{2} - 7x^{3} y^{2}

OpenStudy (mathstudent55):

Terms that have the same variable parts are like terms and can be combined together. For example, 5a - 7a = -2a

OpenStudy (anonymous):

is the answer -2x^3y^2?

OpenStudy (mathstudent55):

correct since the variable parts are the same, x^3y^2, the terms are like terms and can be combined just like you did Good job!

OpenStudy (anonymous):

i have another question

OpenStudy (mathstudent55):

go ahead

OpenStudy (anonymous):

how do i simplify (3m^2+2m-8) + (4m^2-5m+6)

OpenStudy (anonymous):

?

OpenStudy (mathstudent55):

You have two trinomials being added together.

OpenStudy (anonymous):

@mathstudent55 why does your smartscore stick at 99 lol

OpenStudy (mathstudent55):

The first step is to drop both pairs of parentheses. Since there is no minus sign outside to the left of each set of parentheses, the parentheses are unnecessary.

OpenStudy (mathstudent55):

I have no idea how the smart score stuff works.

OpenStudy (anonymous):

would it be 7m^2...

OpenStudy (anonymous):

i dont know what to do next

OpenStudy (mathstudent55):

(3m^2+2m-8) + (4m^2-5m+6) = 3m^2+2m-8 + 4m^2-5m+6 Now you have the same problem without the parentheses. You can only add like terms together. Like terms have exactly the same variable part. 3m^2 and 4m^2 are like terms.

OpenStudy (mathstudent55):

Let me now rearrange it to make it clearer for you. I'l group the pairs of like terms together.

OpenStudy (anonymous):

it would be 7m^2 right?

OpenStudy (mathstudent55):

(3m^2+2m-8) + (4m^2-5m+6) = 3m^2+2m-8 + 4m^2-5m+6 = 3m^2 + 4m^2 + 2m - 5m - 8 + 6

OpenStudy (mathstudent55):

Yes, you start with 7m^2. Then combine the two terms with m, then combine the two terms without variable.

OpenStudy (mathstudent55):

I have to go now. I'll be back in around 15 minutes.

OpenStudy (anonymous):

ok thanks

OpenStudy (mathstudent55):

I'm back. Did you finish it?

OpenStudy (anonymous):

is it 7m^2-3m-14?

OpenStudy (mathstudent55):

Notice you have -8 +6, not +8 +6

OpenStudy (anonymous):

so what would it be?

OpenStudy (mathstudent55):

3m^2 + 4m^2 + 2m - 5m - 8 + 6 = 7m^2 - 3m - 2 You got the m^2 part and the m part correct.

OpenStudy (anonymous):

u teach me more than my math teacher

OpenStudy (mathstudent55):

thanks

OpenStudy (mathstudent55):

Do you have more questions? I can help with for another 15 minutes or so. Then I have to go.

OpenStudy (anonymous):

how do i solve 3n(4n^2+5n?

OpenStudy (mathstudent55):

There is a close parentheses missing. Where does it go?

OpenStudy (mathstudent55):

Is it 3n(4n^2+5n)?

OpenStudy (anonymous):

yes

OpenStudy (mathstudent55):

In this case, you need to use the distributive porperty to distribute 3n by what is inside the parentheses.

OpenStudy (mathstudent55):

Think of this simple example of the distributive property: What is 2(3 + 4) using the distributive property? 2(3 + 4) = 2*3 + 2*4 = 6 + 8 = 14 What is 2(3 + 4) without using the distributive property? 2(3 + 4) = 2(7) = 14 As you can see the result, 14, is the same.

OpenStudy (mathstudent55):

The second method involved 1 addition, 3 + 4 = 7, then one multiplication, 2 * 7 = 14. It is easier to do than to do it with the distributive property.

OpenStudy (mathstudent55):

The distributive property is useful when you can't combine what is inside the parentheses because they are not like terms. Look at this example: What is 2(a + 3) = ?

OpenStudy (mathstudent55):

You can't do a + 3 first since a and 3 are not like terms, so the only way to simplify 2(a + 3) is by using the distributive property. Using the distributive property, you get 2(a + 3) = 2*a + 2*3 = 2a + 6 That is as simplified as you can get since 2a and 6 are not like temrs and can't be added together.

OpenStudy (mathstudent55):

Now let's get back to our problem: 3n(4n^2+5n) The terms inside the parentheses are not "like terms" since n^2 and n are not the same. The terms inside the parentheses cannot be added together. The only way to simplify this expression is by using the distributive property. 3n(4n^2+5n) = 3n * 4n^2 + 3n * 5n = 12n^3 + 15n^2

OpenStudy (mathstudent55):

You haven't answered in a while. I hope I haven't put you to sleep. Are you following this?

OpenStudy (anonymous):

haha sorry my computer died

OpenStudy (mathstudent55):

Take a while to read what I wrote and ask any questions you may have.

OpenStudy (anonymous):

i cant combine 12n^3 and 15n^2 because the exponents are different right?

OpenStudy (mathstudent55):

correct

OpenStudy (anonymous):

how would i solve 4k^2(3-4k)

OpenStudy (anonymous):

?

OpenStudy (mathstudent55):

Once again, since you can't combine the terms inside the parentheses bec thay are not like terms, you must use the distributive property. 4k^2(3-4k) = 4k * 3 - 4k * 4k = 12k - 16k^2

OpenStudy (anonymous):

how do i solve 18s-63?

OpenStudy (mathstudent55):

What exactly are the instructions with this problem?

OpenStudy (mathstudent55):

You can't combine together 18s and - 63 because they are not like terms.